Find cv and cp for a certain fluid if the properties are related as follows: u = (196 + 0.718t) KJ/kgm and pv = [0.287(toc + 273)] where v is in m3/kgm and p is in KN/m2.


1
Expert's answer
2021-01-25T04:09:46-0500

U = (196 + 0.718t) KJ/kg

Cv="\frac{dU}{dT}=\frac{d(196+0.718t)}{dT}"


"C_v=0.718 \frac{kJ}{kgK}"

PV=[0.287(toc + 273)] 

"C_p=C_v+P(\frac{\delta V}{dT})"


"C_p=0.718+.287=1.005 \frac{Kj}{kgK}"